if the sides of a square are increased by 3cm its' area becomes 100 square cm. find the length of the sides of the original square.
Answer by checkley77(12844)

(X+3)^2=100
X^2+6X+9=100
X62+6X+9-100=0
X^2+6X-91=0
(X-7)(X+13)=0
X-7=0
X=7cm. FOR THE ORIGINAL SIDE OF THE SQUARE.